# Ratio Covers of Convex Sets and Optimal Mixture Density Estimation

### v1

@@ Abstract
Our analysis of the mixture case hinges on two new covering results. First, we provide a sharp, distribution-free upper bound on the local Hellinger entropy of the class of mixtures of $M$ distributions. Second, we prove an optimal ratio covering theorem for convex sets: for every convex compact set $K\subset\mathbb{R}_{+}^{d}$, there exists a subset $A\subset K$ with at most $2^{O(d)}$ elements such that each element of $K$ is coordinate-wise dominated by an element of $A$ up to a universal constant factor. This geometric result is of independent interest; notably, it yields new cardinality estimates for $\varepsilon$-approximate Pareto sets in multi-objective optimization when the feasible objective set is convex.
@@ 1 Introduction
We revisit two classical distribution estimation problems. Let $\mathcal{X}$ be a measurable space endowed with a measure $\mu$. In addition, let $p_{1},\ldots,p_{M}$ be a collection of $M\geqslant 2$ known densities on $\mathcal{X}$ with respect to $\mu$, with no assumptions on these densities (in particular, the corresponding distributions need not satisfy any moment or tail conditions, and their supports may be different). We observe an i.i.d. sample $X_{1},\ldots,X_{n}$ drawn from some unknown density $p^{\star}$ for which we only assume either
$p^{\star}\in\{p_{1},\ldots,p_{M}\}\ \,\textrm{(Model Aggregation),\quad or}\quad p^{\star}\in\operatorname{conv}\{p_{1},\ldots,p_{M}\}\ \,\textrm{(Convex Aggregation).}$
@@
Even in the Model Aggregation setup, this task is known to be harder than standard hypothesis testing. Indeed, it is easy to construct a pair of densities $p,q$ such that no test can reliably distinguish them, while $\operatorname{KL}(p,q)=\operatorname{KL}({q},{p})=+\infty$. This makes the naive approach of using hypothesis tests to choose between candidates risky, and suggests that even in Model Aggregation the estimator should output a mixture density, namely $\widehat{p}\in\operatorname{conv}\{p_{1},\ldots,p_{M}\}$.
In Model Aggregation, the state-of-the-art result is due to the seminal work of Yang and Barron Yang and Barron (1999); Yang (2000). Their estimator (the progressive mixture rule) traces back to Barron (1987) and an independent line of work by Catoni Catoni (1997, 2004). For $M$ densities they construct an estimator $\widetilde{p}=\widetilde{p}(X_{1},\ldots,X_{n})$ such that
@@
As we discuss below, the estimator of Yang and Barron provably does not yield the optimal high-probability bound in the Model Aggregation setup, and neither does Bayesian model averaging. For Convex Aggregation, Theorem 2 of Mourtada (2025) shows that any estimator not using $\delta$ must incur suboptimal high-probability bounds, ruling out many classical estimators. Perhaps more surprisingly, for Convex Aggregation even the optimal in-expectation bound is unknown in full generality. The best known general upper bound follows from regret bounds over the simplex Hazan et al. (2007) together with an online-to-batch conversion (related log-loss regret for the simplex goes back to Krichevsky–Trofimov Krichevsky and Trofimov (1981) and to the universal portfolio analysis of Cover Cover (1991)). This standard route leads to averaging the per-round Bayes predictors, as in the analysis of Yang and Barron. Equivalently, running exponential weights over the simplex yields, in the Convex Aggregation regime,
@@
This bound is known to be suboptimal at least in the canonical special case where the $M$ distributions correspond to Dirac masses at different elements, namely discrete distribution estimation. In particular, over the years a series of papers Bhattacharyya et al. (2021); Catoni (1997); Forster and Warmuth (2002); Braess and Sauer (2004); Mourtada and Gaïffas (2022) showed that the optimal risk in this case scales as $\frac{M}{n}$ (up to constants), without the additional multiplicative $\log\left(\frac{n}{M}\right)$ factor in (2), often via analyses of the Laplace estimator. Moreover, there has been strong recent interest in high-probability bounds for the discrete case Canonne et al. (2023); van der Hoeven et al. (2023); Han et al. (2023); van der Hoeven et al. (2025), leading finally to optimal high-probability guarantees Mourtada (2025) for discrete distributions. The techniques used in these works are largely specific to the discrete setting and differ from those developed here.
@@ 1.1 Our contributions
Our main statistical contributions are optimal high-probability guarantees for both model aggregation and convex aggregation (as well as optimal in-expectation rates for convex aggregation), for arbitrary dictionary of densities. Here and in what follows $a\lesssim b$ means that there exists an absolute constant $c>0$ such that $a\leqslant cb$.
For Model Aggregation we present an estimator $\widehat{p}_{\operatorname{MA}}$, and for Convex Aggregation an estimator $\widehat{p}_{\operatorname{CA}}$, such that with probability at least $1-\delta$,
$\operatorname{KL}\big(p^{\star},\widehat{p}_{\operatorname{MA}}\big)\lesssim\frac{\log M\log(1/\delta)}{n},\qquad\text{and}\qquad\operatorname{KL}\big(p^{\star},\widehat{p}_{\operatorname{CA}}\big)\lesssim\frac{M+\log M\log(1/\delta)}{n}.$
These guarantees are stated in Theorem˜2 and Theorem˜6. Both rates are unimprovable in general, up to multiplicative constant factors.
@@
A standard approach to these problems is to use metric entropy estimates: one derives statistical guarantees by analyzing coverings of the relevant space of distributions. As mentioned above, existing implementations of this strategy have not yielded sharp guarantees in our setting. In this work, we develop tools that make a metric-entropy approach sharp. There are two main obstacles: (i) we do not have sharp covering numbers for mixture models in KL divergence, and (ii) even with sharp covering numbers, the usual reductions would still not lead to sharp statistical bounds.
@@
Our approach relies on three ingredients. First, for both problems we construct a preliminary estimator that is close to $p^{\star}$ in Hellinger distance. This metric is more amenable to entropy methods, but we need stronger local covering bounds than were previously available (to the best of our knowledge, this remained open; see Baraud and Birgé (2018) for the state of the art). We establish the required entropy estimate (Theorem 3) in Section 3. Second, we construct an $\varepsilon$-cover in KL divergence within the resulting local Hellinger ball. Third, we treat this KL cover as a finite list of candidates and feed it into the recent estimator of van der Hoeven et al. (2023) under bounded density ratios, which returns a distribution whose KL risk is nearly as good as the best candidate in the list. We briefly gloss over some additional technicalities. In particular, we mix candidates in a specific way to satisfy the conditions required by the estimator. For completeness, we provide a self-contained (less general) version of this estimator in Appendix˜A.
The second ingredient, namely constructing a KL cover of a local Hellinger ball, is a key result of the paper. For reasons explained later, it reduces to the following more general geometric problem. We want a small subset $A$ of the Hellinger ball such that for any coefficient vector $\theta=(\theta_{1},\dots,\theta_{M})$ in the ball, there exists $\phi=(\phi_{1},\dots,\phi_{M})\in A$ whose coordinates are not much smaller than those of $\theta$ (in terms of their ratios). This motivates the notion of ratio covers:
Definition 1.
In Section 4, we provide two different proofs that convex sets admit small ratio covers. The next geometric result is our second main contribution.
Theorem 1.
@@ 1.2 Further applications of ratio covers
Let $K\subseteq\mathbb{R}_{+}^{d}$ represent attainable values in a multi-objective optimization problem. For instance, a point $x=(x_{1},\dots,x_{d})\in K$ could represent producing $x_{1}$ units of product 1, $x_{2}$ units of product 2, and so on. In general, the Pareto-optimal set may be infinite. The influential work of Papadimitriou and Yannakakis Papadimitriou and Yannakakis (2000) shows that there exists a small subset of $K$ that approximately dominates all points in $K$. In our terminology, a set that $(1+\varepsilon)$-dominates is exactly a $(1+\varepsilon)$-ratio cover. Their bounds for general $K$ depend on $d,\varepsilon$ and on the range parameter $R=\max_{i\in[d]}\frac{\max_{x\in K}x_{i}}{\min_{y\in K}y_{i}}$, where here and in what follows $d]=\{1,\ldots,d\}$. In Section 4.2, we use our ratio cover machinery (under convexity of $K$) to remove the dependence on $R$. We expect this to be useful in settings where $K$ is convex, for example when $K$ is the feasible region of a linear program.
@@
As remarked earlier, Yang and Barron provide a general recipe for converting KL covering numbers into in-expectation error bounds. While very general, their method often introduces an unnecessary logarithmic factor. In Section 6.1, we sharpen the Yang–Barron guarantee in settings where the class is convex and amenable to our ratio-cover approach. The plan is: (i) obtain an estimator with squared Hellinger error $\varepsilon$ (often via entropy methods), (ii) use ratio covers to build a KL cover of the local Hellinger ball, and (iii) run the Yang–Barron estimator using this local KL cover. The key improvement is that we only need a KL cover locally, and the ratio-cover structure remains helpful even when the initial Hellinger estimator occasionally lands outside the target $\varepsilon$-ball. This yields our in-expectation guarantee for convex aggregation, scaling as $\frac{M}{n}$ up to constants.
@@ 1.4 Notation and preliminary results
Throughout this work, $\log(\cdot)$ denotes the natural logarithm. Let $[k]$ denote $\{1,2,\ldots,k\}$. $\widehat{L}_{n}$ is the normalized negative log-likelihood; namely, $\widehat{L}_{n}(p)=-\frac{1}{n}\sum\limits_{i=1}^{n}\log(p(X_{i}))$. Squared Hellinger divergence is denoted as $\operatorname{H}^{2}(f,g)=\frac{1}{2}\int_{\mathcal{X}}(\sqrt{f(x)}-\sqrt{g(x)})^{2}\mathrm{d}\mu(x)$. We will use the following relation between KL divergence and the Hellinger distance.
@@
Assume that $f,g$ are two densities on $\mathcal{X}$ with respect to $\mu$. It holds that
$\operatorname{KL}(f,g)\leqslant\frac{2}{(\sqrt{e}-1)^{2}}\,\operatorname{H}^{2}(f,g)\,\max\left\{1,\ \sup_{x\in\mathcal{X}}\log\left(\frac{f(x)}{g(x)}\right)\right\}.$
@@
Exact construction of the estimator as well as the self-contained proof of Lemma˜2 are presented in Appendix˜A. As a short comment, the estimator $\widehat{p}$ is a specific variance-adjusted version of the standard progressive mixture rule. Note that Lemma˜2 does not solve the model aggregation problem introduced above as it requires that the density ratios are bounded, an assumption we aim to avoid in this paper.

### v2

@@ Abstract
Our analysis of the mixture case hinges on two new covering results. First, we provide a sharp, distribution-free upper bound on the local Hellinger entropy of the class of mixtures of $M$ distributions. Second, we prove an optimal ratio covering theorem for convex sets: for every convex compact set $K\subset\mathbb{R}_{+}^{d}$, there exists a subset $A\subset K$ with at most $2^{O(d)}$ elements such that each element of $K$ is coordinate-wise dominated by an element of $A$ up to a universal constant factor. This geometric result is of independent interest; notably, it yields new cardinality estimates for $\varepsilon$-approximate Pareto sets in multi-objective optimization with convex feasible set.
@@ 1 Introduction
We revisit two classical distribution estimation problems. Let $\mathcal{X}$ be a measurable space endowed with a measure $\mu$. In addition, let $p_{1},\ldots,p_{M}$ be a collection of $M\geqslant 2$ known densities on $\mathcal{X}$ with respect to $\mu$, with no assumptions on these densities (in particular, the corresponding distributions need not satisfy any moment or tail conditions, and their supports may differ). We observe an i.i.d. sample $X_{1},\ldots,X_{n}$ drawn from some unknown density $p^{\star}$ for which we only assume either
  1. 1.
    $p^{\star}\in\{p_{1},\ldots,p_{M}\}$, in the Model Aggregation setting;
  2. 2.
    or $p^{\star}\in\operatorname{conv}\{p_{1},\ldots,p_{M}\}$, in the Convex Aggregation (or mixture density estimation) setting.
@@
In this work, our primary aim is to characterize the best possible high-probability guarantees for both Model Aggregation and Convex Aggregation. More precisely, given $M,n\geqslant 2$ and $\delta\in(0,1/2)$, we aim to identify (up to absolute constant factors) the smallest quantities $\psi_{\mathrm{MA}}(M,n,\delta)$ and $\psi_{\mathrm{CA}}(M,n,\delta)$ such that, for every probability densities $p_{1},\dots,p_{M}$ on a measure space $(\mathcal{X},\mu)$:
  1. 1.
    there exists an estimator $\widehat{p}_{\mathrm{MA}}$ such that, for every $p^{\star}\in\{p_{1},\dots,p_{M}\}$, with probability at least $1-\delta$ over $X_{1},\dots,X_{n}$ i.i.d. from $p^{\star}$ one has
    $\operatorname{KL}({p^{\star}},{\widehat{p}_{\mathrm{MA}}})\leqslant\psi_{\mathrm{MA}}(M,n,\delta);$ (1)
  2. 2.
    there exists an estimator $\widehat{p}_{\mathrm{CA}}$ such that, for every $p^{\star}\in\mathop{\mathrm{conv}}\{p_{1},\dots,p_{M}\}$, with probability at least $1-\delta$ over $X_{1},\dots,X_{n}$ i.i.d. from $p^{\star}$ one has
    $\operatorname{KL}({p^{\star}},{\widehat{p}_{\mathrm{CA}}})\leqslant\psi_{\mathrm{CA}}(M,n,\delta).$ (2)
The fact that we make no assumption on the densities [^1] Given probability distributions $P_{1},\dots,P_{M}$ on $\mathcal{X}$, the existence of a dominating measure $\mu$ such that each $P_{j}$ admits a density $p_{j}$ with respect to $\mu$ is not restrictive, as one may simply take $\mu=P_{1}+\dots+P_{M}$. $p_{1},\dots,p_{M}$ induces some difficulties even for Model Aggregation. Indeed, in this case it is straightforward to show (see Fact 1 in Appendix B) that no selection method $\widehat{p}$ returning an element of the dictionary $\{p_{1},\dots,p_{M}\}$ achieves any meaningful distribution-free guarantee. In particular, this rules out arguably the simplest method, namely the maximum likelihood estimator (MLE). For this reason, the procedures we will consider will return mixtures of distributions from the dictionary, even for Model Aggregation.
Aside from the MLE, the most common method for Model Aggregation is Bayesian model averaging, which returns a mixture of the densities $p_{1},\dots,p_{M}$ weighted by their Bayesian posterior weights (say, starting from a uniform prior distribution over $p_{1},\dots,p_{M}$). In Appendix B, we show that this standard method is also suboptimal in the high-probability regime. Specifically, we obtain matching upper and lower bound on the tail performance of Bayes model averaging in the case where $M=2$, which show that this method achieves nontrivial yet suboptimal guarantees.
We note that there is a rich literature on Kullback-Leibler density estimation, both in the Model aggregation Barron (1987); Yang and Barron (1999); Yang (2000); Catoni (1997); Juditsky et al. (2008); Audibert (2007, 2009); Rigollet (2012); Butucea et al. (2017); Baraud (2021); van der Hoeven et al. (2023) and Convex aggregation Li and Barron (1999); Li (1999); Rakhlin et al. (2005); Dalalyan and Sebbar (2018) settings. However, our point of view differs from that of the aggregation literature. Indeed, a key emphasis in these works is the effect of model misspecification, where $p^{\star}$ does not belong to $\{p_{1},\dots,p_{M}\}$; while—with two exceptions discussed below—the ratios of densities $p_{1},\dots,p_{M}$ are assumed to be bounded, with the guarantees diverging when the bound on density ratios diverges. In contrast, we consider arbitrary densities $p_{1},\dots,p_{M}$, for which optimal high-probability guarantees are unknown even in the well-specified case (where $p^{\star}$ satisfies either of the two assumptions above). For these reasons, we focus in this work on optimal guarantees in the well-specified case, although we provide in Section 7 extensions to the misspecified case with additional logarithmic factors.
An important special case of mixture density estimation in which the boundedness condition fails to hold is that of estimation of discrete distributions, where one aims to estimate an arbitrary distribution on a finite set (say, $\{1,\dots,M\}$). This can be framed as mixture density estimation, where $\mathcal{X}=\{1,\dots,M\}$, $\mu$ is the counting measure (i.e., $\mu(A)$ is the cardinality of $A$ for $A\subset\mathcal{X}$), and for $1\leqslant j\leqslant M$ the density $p_{j}:\mathcal{X}\to\mathbb{R}^{+}$ is defined by $p_{j}(k)=1$ if $k=j$ and $p_{j}(k)=0$ if $k\neq j$. (In other words, any distribution on $\mathcal{X}$ is a mixture of Dirac masses.) In this case, it is relatively straightforward to show, and has been known for a long time Catoni (1997); Forster and Warmuth (2002); Braess and Sauer (2004); Mourtada and Gaïffas (2022), that optimal in-expectation guarantees are of order $M/n$. In contrast, the study of high-probability bounds is markedly more delicate and has received significant interest in recent years Bhattacharyya et al. (2021); Han et al. (2023); Canonne et al. (2023); van der Hoeven et al. (2023); Mourtada (2025); van der Hoeven et al. (2025), leading finally to optimal guarantees in Mourtada (2025). However, the techniques used in these works are largely specific to the discrete setting, and it is unclear how to extend them to the general mixture setting.
We now discuss existing guarantees in Model Aggregation and Convex Aggregation that apply to arbitrary densities $p_{1},\dots,p_{M}$—both of which only hold in expectation.
For Model Aggregation, optimal guarantees in expectation have been obtained in seminal work of Yang and Barron Yang and Barron (1999); Yang (2000) and independently by Catoni Catoni (1997, 2004). The corresponding estimator $\widetilde{p}$ (the progressive mixture rule) traces back to Barron (1987) and consists in a version of Bayes model averaging with additional averaging over sample sizes. For $M$ densities, it achieves the following guarantee:
@@
While this conclusively addresses the important question of optimal in-expectation guarantees, it leaves the more delicate question of optimal high-probability guarantees unanswered. Specifically, the upper bound (3) implies that $\psi_{\mathrm{MA}}(M,n,\delta)\leqslant\log(M)/(n\delta)$, but this bound is highly suboptimal in the high-confidence regime of small $\delta$. In fact, as we discuss in Appendix B, the progressive mixture rule achieves suboptimal high-probability guarantees, hence another estimator $\widehat{p}$ is required.
For Convex Aggregation, optimal guarantees are unknown even in expectation. The best known general upper bound is achieved by a continuous version of the progressive mixture rule $\widetilde{p}$, applied to the probability simplex with uniform prior distribution. Combining regret bounds for the Bayes mixture method Hazan et al. (2007) (related log-loss regret guarantees for the simplex go back to Krichevsky–Trofimov Krichevsky and Trofimov (1981) and to the universal portfolio analysis of Cover Cover (1991), see also Jézéquel et al. (2025) for recent progress on this problem) with an online-to-batch conversion gives the following guarantee:
@@
This bound is known to be suboptimal at least in the canonical special case of discrete distribution estimation discussed above, due to the presence of the $\log(1+n/M)$ factor. In addition, as in the Model Aggregation case, it does not address the question of optimal high-probability guarantees.
@@ 1.1 Our contributions
Our main statistical contributions are optimal high-probability guarantees for both Model Aggregation and Convex Aggregation (as well as optimal in-expectation rates for Convex Aggregation), for arbitrary dictionary of densities. In what follows $a\lesssim b$ means that there exists an absolute constant $c>0$ such that $a\leqslant cb$.
  1. 1.
    Theorem 1: for any $M,n\geqslant 2$ and $\delta\in(e^{-n},e^{-1})$, the optimal high-probability rate (1) of Model Aggregation scales as
    $\psi_{\mathrm{MA}}(M,n,\delta)\asymp\frac{\log M\log(1/\delta)}{n}.$
  2. 2.
    Theorem 6: for any $n\geqslant M\geqslant 2$ and $\delta\in(e^{-n},e^{-1})$, the optimal high-probability rate (2) of Convex Aggregation scales as
    $\psi_{\mathrm{CA}}(M,n,\delta)\asymp\frac{M+\log M\log(1/\delta)}{n}.$
Specifically, the lower bounds above have been established in the special case of discrete distributions Mourtada (2025); van der Hoeven et al. (2025); our contribution is to establish matching upper bounds for arbitrary dictionaries.
@@
A standard approach to these problems is to use metric entropy estimates: one derives statistical guarantees by analyzing coverings of the relevant class of distributions. As mentioned above, existing implementations of this strategy have not yielded sharp guarantees in our setting. In this work, we develop tools that make a metric-entropy approach sharp. There are two main obstacles: (i) we do not have sharp covering numbers for mixture models in KL divergence, and (ii) even with sharp covering numbers, the usual reductions would still not lead to sharp statistical bounds.
@@
Our approach relies on three ingredients. First, for both problems we construct a preliminary estimator that is close to $p^{\star}$ in Hellinger distance. This metric is more amenable to entropy methods, but for the mixture case we need stronger local covering bounds than were previously available (to the best of our knowledge, this remained open; see Baraud and Birgé (2018) for the state of the art). We establish the required entropy estimate (Theorem 2) in Section 3. Second, we construct an $\varepsilon$-cover in KL divergence within the resulting local Hellinger ball. Third, we treat this KL cover as a finite set of candidates and feed it into the recent estimator of van der Hoeven et al. (2023) under bounded density ratios, which returns a distribution whose KL risk is nearly as good as the best candidate in the set. We briefly gloss over some additional technicalities. In particular, we mix candidates in a specific way to satisfy the conditions required by the estimator. For completeness, we provide a self-contained (less general) version of this estimator in Appendix˜A.
The second ingredient, namely constructing a KL cover of a local Hellinger ball, is a critical step of the proof. For reasons explained later, it reduces to the following more general geometric problem. Given a convex compact subset $K\subset\mathbb{R}_{+}^{d}$, we would like to find a finite subset $A\subset K$ with controlled cardinality satisfying the following (one-sided) multiplicative approximation guarantee: for every $\theta=(\theta_{j})_{1\leqslant j\leqslant d}\in K$, there exists an element $\phi=(\phi_{j})_{1\leqslant j\leqslant d}\in A$ such that $\theta_{j}\leqslant c\,\phi_{j}$ for every $j=1,\dots,d$, where $c>1$ is an absolute constant. We call such a set $A$ a $c$-ratio cover of $K$ (Definition 1).
In Section 4, we show that convex sets admit small ratio covers. In particular, we establish the following optimal cardinality estimate for ratio covers of compact convex sets, which is one of the main contributions of this work. We expect this covering result to be of independent interest, beyond its application to estimation of mixture distributions.
  1. 3.
    We show that there exist universal constants $C_{1},C_{2}>1$ such that the following holds: for every $d\geqslant 1$, every convex compact subset of $\mathbb{R}_{+}^{d}$ admits a $C_{1}$-ratio cover with at most $C_{2}^{d}$ elements (Theorem˜3).
@@ 1.2 Further applications of ratio covers
Let $K\subset\mathbb{R}_{+}^{d}$ represent the set of attainable values in a multi-objective optimization problem. For instance, a point $x=(x_{1},\dots,x_{d})\in K$ could represent producing $x_{1}$ units of product 1, $x_{2}$ units of product 2, and so on. In general, the Pareto-optimal set may be infinite. The influential work of Papadimitriou and Yannakakis Papadimitriou and Yannakakis (2000) shows that there exists a small subset of $K$ that approximately dominates all points in $K$. In our terminology, a set that $(1+\varepsilon)$-dominates $K$ is exactly a $(1+\varepsilon)$-ratio cover. Their bounds for general $K$ depend on $d,\varepsilon$ and on the range parameter $R=\max_{i\in[d]}\frac{\max_{x\in K}x_{i}}{\min_{y\in K}y_{i}}$, where here and in what follows $d]=\{1,\ldots,d\}$. In Section 4.2, we use our ratio cover machinery (under convexity of $K$) to remove the dependence on $R$. We expect this to be useful in settings where $K$ is convex, for example when $K$ is the feasible region of a linear program.
@@
As remarked earlier, Yang and Barron provide a general recipe for converting KL covering numbers into in-expectation error bounds. While very general, their method often introduces an unnecessary logarithmic factor. In Section 6.1, we sharpen the Yang–Barron guarantee in settings where the class is convex and amenable to our ratio-cover approach. The plan is: (i) obtain an estimator with squared Hellinger error $\varepsilon$ (often via entropy methods), (ii) use ratio covers to build a KL cover of the local Hellinger ball, and (iii) run the Yang–Barron estimator using this local KL cover. The key improvement is that we only need a KL cover locally, and the ratio-cover structure remains helpful even when the initial Hellinger estimator occasionally lands outside the target $\varepsilon$-ball (thanks to Lemma˜5). This yields our in-expectation guarantee for convex aggregation, scaling as M/n{M}/{n} up to constants.
@@ 1.4 Notation and preliminary results
Let $[k]$ denote $\{1,2,\ldots,k\}$. $\widehat{L}$ is the normalized negative log-likelihood, namely
$\widehat{L}(p)=-\frac{1}{n}\sum\limits_{i=1}^{n}\log p(X_{i}).$ (5)
Squared Hellinger divergence is denoted as
$\operatorname{H}^{2}(p,q)=\frac{1}{2}\int_{\mathcal{X}}\big(\sqrt{p(x)}-\sqrt{q(x)}\big)^{2}\mathrm{d}\mu(x).$ (6)
We will use the following relation between KL divergence and the Hellinger distance.
@@
Assume that $p,q$ are two densities on $\mathcal{X}$ with respect to $\mu$. It holds that
$\operatorname{KL}(p,q)\leqslant\frac{2}{(\sqrt{e}-1)^{2}}\,\operatorname{H}^{2}(p,q)\,\max\left\{1,\ \sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\right\}.$
@@
Exact construction of the estimator as well as the self-contained proof of Lemma˜2 are presented in Appendix˜A. As a short comment, the estimator $\widehat{p}$ is a specific variance-adjusted version of the standard progressive mixture rule. Note that Lemma˜2 does not solve the model aggregation problem introduced above as it requires that the density ratios are bounded, an assumption we aim to avoid in this work.